# An egonet-based approach to effective weighted network comparison

### v2

@@ Abstract
In the literature, a baseline method for defining the dissimilarity between two weighted graphs is to calculate the difference between one or more global (scalar) quantities, such as the clustering coefficient, average shortest path, or diameter. This approach, however, has been proven to have limited success on binary networks [21, 29] and there is no reason to believe that it should be more effective in an even more challenging context. For benchmark purposes, we will evaluate the performance of using the global (weighted) clustering coefficient as a dissimilarity metric. In the same spirit we will also compare our proposal to spectral metrics, which are based on the comparison of the set of eigenvalues of the (weighted) adjacency or Laplacian matrices[33]. In the set of benchmark methods we will also evaluate Portrait Divergence, which defines a distance by comparing the ”portrait” of the two graphs, i.e., a matrix that encodes the distribution of the lengths of the shortest-path[34], and behaves very well in classifying binary networks[29]. It has been generalized to define a measure of dissimilarity between weighted networks (Ref.[34], Supplemental Material), although it has not been extensively tested in this regard. Finally, we note that the class of graphlet-based alignment-free methods [31, 32], which guarantees very high-level performance on binary networks [29, 30], does not yet have a generalization to weighted networks, as discussed in Ref.[35].
@@
The classification capabilities of the proposed ”weighted ego-distances” will be extensively evaluated on the testbed defined above. More precisely, the aim is to discriminate pairs of graphs originating from the same model from those originating from different models, with size and density (i.e., number of nodes and edges) acting as confounding factors. Note that the definition of ”model” includes the weighting scheme, so, for example, a dissimilarity measure should be able to recognize that two Barabási-Albert networks have different weighting schemes, even if they have exactly the same topology. The results of the experiments show that the proposed weighted ego-distances are able to outperform all distances used for comparison, and therefore should be considered, to date, the most advanced method for comparing undirected weighted networks.
The paper concludes with two application examples. In the first, a weighted ego-distance is used to compare the effects of different algorithms to filter weighted networks, i.e., to derive a ”backbone” by eliminating a large number of insignificant edges while maintaining most of the total network weight[38]. Three filtering schemes are evaluated, namely the trivial hard threshold, the classical disparity filter[39], and the more recent Pólya filter[40]. Their application to two different datasets (the world trade network and the US airport network) reveals the superiority of the Pólya filter, in the sense that, at a given level of pruning, it is able to preserve the egonet (i.e., local) features of the original graph to a greater extent. The second application example concerns a temporal sequence of correlation networks, each obtained by analyzing the relationships between 3-month long time series referring to the stocks that define the S&P100 index. The pairwise comparison between the time-indexed correlation networks confirms a well-known result, namely that, given the greater coherence between the time series, the correlation graphs associated with financial crises have significantly different characteristics from those characterizing ordinary periods[41, 17, 42].
@@ Results
We evaluate the effectiveness of the ego-distances proposed above by generalizing the experimental setup defined in Refs. [31, 32, 30] with reference to unweighted networks. However, it is necessary to extend the above framework to the case of weighted networks for which, to the best of our knowledge, a standard benchmark pool of network models to be used for classification has never been defined to date.
@@
  • ER-x: Erdős-Rényi model with Uniform (ER-U), Random (ER-R), or Degree-dependent (ER-D) weighting
  • BA-x: Barabási-Albert model with Uniform (BA-U), Random (BA-R), or Degree-dependent (BA-D) weighting
  • GEO-x: geometric random graph model with Uniform (GEO-U), Random (GEO-R), or Degree-dependent (GEO-D) weighting
  • YJBT: Yook-Jeong-Barabási-Tu model[36]
  • AK-x: Antal-Krapisvky model with Random (AK-R) or Exponential (AK-E) weighting[37]
@@
Figure 3: Example distributions of (normalized) weighted degree (left column), clustering coefficient (middle column), and egonet persistence (right column) for the $12$ synthetic network models ($N=1000$, $\rho=0.01$). Degree-homogeneous networks are in the top row, degree-heterogeneous networks in the bottom row.
@@
For each pair of networks, we calculate the eight ego-distances defined above (see section Ego-distances). To have a comparison in terms of classification performance, we also calculate the following benchmark distances[29] (see section Methods for details):
@@
  • $D_{PDiv}$ (Portrait Divergence): It is based on the comparison of the portrait matrices of the two networks, which encode the distribution of the shortest-path lengths of the graphs[34].
@@
Table 1 summarizes, in terms of AUPR value, the results of the classification experiments described above. To combine the performance obtained by comparing graphs with different and with the same size/density, we also included the average of the two AUPR values (last column) as an overall indicator. In general, all ego-distances that use a combination of egonet features (i.e., $D_{SUM}$, $D_{cp}$, $D_{dc}$, $D_{dp}$, and $D_{dcp}$) perform better than the benchmark distances ($D_{C_{global}},D_{SP-W/L},D_{PDiv}$) in most cases. A partial exception is $D_{PDiv}$ (Portrait Divergence), which performs similarly to ego-distances when comparing graphs of the same size/density, although performance drops significantly more than ego-distances when considering mixed size/density. Overall, $D_{dcp}$, which exploits all the information available in the most extensive form, obtains the best performance on average. A viable and computationally simpler alternative is $D_{SUM}$, which achieves only slightly worse performance but is computationally lighter. The Precision-Recall curves of Fig. 4 (left panel) show that the ego-distances which exploit all the available information ($D_{dcp}$ and $D_{SUM}$) achieve similar behavior to each other, and outperform the benchmark distances ($D_{PDiv}$ and to a larger extent $D_{C_{global}}$). $D_{dcp}$ provides the best trade-off at high Recall values, as evidenced by the largest F1 value slightly above $0.5$ (we recall that F1 is the harmonic mean of Precision and Recall). On the other hand, when the comparison is limited to networks with the same size/density, Fig. 4 (right panel) confirms that, as underlined in Table 1, $D_{dcp}$, $D_{SUM}$, and $D_{PDiv}$ achieve similar performance, while $D_{C_{global}}$ follow with a large gap.
@@
distance all sizes-densities same size-density avg
@@
Figure 4: Precision-Recall curves obtained by classifying the complete set of networks (12 models $\times$ 3 sizes $\times$ 3 densities $\times$ 10 replications = 1080 networks) with two of the ego-distances defined in the section Methods ($D_{SUM}$, $D_{dcp}$) and two of the benchmark distances ($D_{C_{global}}$, $D_{PDiv}$) (the curves related to all other distances are omitted for readability). For each Recall-Precision point, the F1 value (i.e., the harmonic mean of Precision and Recall) is specified by the dotted contour line. Left: Performance is assessed by mixing sizes/densities; Right: Only networks with the same size/density are compared.
@@
Table 2: AUPR (Area Under the Precision/Recall curve) value for the classification of network models, for the ego-distances (top part of the table) and for the benchmark distances (bottom part) (we omit the distances $D_{d}$, $D_{c}$, $D_{p}$, and $D_{SP-W/L}$, whose performance is lower in all cases). The classification is performed separately for networks with homogeneous or heterogeneous degree distribution.
@@
We now summarize the results reported above, in order to provide guidance on which of the proposed ego-distances to adopt. We first observe that our results depend on the experimental dataset, which is rich and diverse but, obviously, does not include all possible network structures: any conclusion must therefore be evaluated with caution. In general, the AUPR (Table 1) should be the primary selection criterion, as it represents the average precision across all possible recall values. Not surprisingly, AUPR highlights the superiority of measures that exploit all available information: $D_{dcp}$ and $D_{SUM}$ have the best (average) performances, and the second is a valid alternative to the first, with lower computational load (see Methods).
@@
The results are summarized in Fig. 5. As expected, for both networks and all filters used, the distance increases monotonically as more and more weight is removed. However, for any given value of the removed weight, the distance from the Hard threshold graph is greater than the distance from the graphs obtained by using the nontrivial filters (Disparity and Pólya), denoting that the former tends to perturb the original features to a larger extent. Furthermore, the Pólya filter produces backbone graphs much more similar to the original one than the Disparity filter, when the weight reduction is small (the two methods approximately coincide when the Pólya filter parameter reaches $a=1$, as theoretically discussed[40]). We note that, at the weight reduction level $R^{f}$ where the two curves become almost coincident, more than 80% of links are removed in both study cases, while maintaining at least 90% of the total weight. Therefore, the greater similarity with the original graph obtained from the Pólya filter is particularly appreciable, since it denotes that, for the same pruning effort, this scheme is able to preserve the egonet features of the original graph to a greater extent.
@@
Let us now use ego-distances to analyze data describing the stocks defining the S&P100 index, which includes most of the largest US companies. We start from the daily closing values $y_{i}^{\tau}$ of all stocks $i$ for the 10-year period January 2007 – December 2016 and use them to calculate the daily returns $y_{i}^{\tau+1}/y_{i}^{\tau}$, which are then standardized to obtain time series $u_{i}^{\tau}$ with zero mean and unit variance. The 10-year time span is then divided into quarters $t=1,2,\ldots,40$ (that is, January 2006-March 2006, April 2006-June 2006, etc.). For each quarter, a correlation network is defined whose $100$ nodes correspond to the stocks. The nodes $(i,j)$ are connected with a weight given by $w_{ij}^{t}=(\rho_{ij}^{t}+1)/2$, where $\rho_{ij}^{t}$ is the Pearson correlation between the daily time series $(u_{i}^{\tau},u_{j}^{\tau})$ for the set of days $\tau$ of the quarter $t$. Thus $w_{ij}^{t}$ ranges from 00 (perfect anti-correlation) to $1$ (perfect correlation), with larger values denoting stronger agreement.
@@ Discussion
A significant limitation to the widespread use of the ego-distances could be their computational cost, which scales unfavorably with network size (see Methods) and therefore limits their applicability to small/medium sized graphs. It should be noted, however, that in principle the computation of egonet features can be fully parallelized, potentially allowing for a considerable reduction in computational requirements. Alternatively (or in combination), one could infer egonet distributions from only a sample of nodes, rather than from the entire network. This last approach, potentially promising for the analysis of large-scale networks, requires the use of appropriate graph sampling techniques[47]. We leave the in-depth evaluation of the effect of sampling on the performance of ego-distances to future research.

### v3

@@ Abstract
In the literature, a baseline method for defining the dissimilarity between two weighted graphs is to calculate the difference between one or more global (scalar) quantities, such as the clustering coefficient, average shortest path, or diameter. This approach, however, has been proven to have limited success on binary networks [21, 29] and there is no reason to believe that it should be more effective in an even more challenging context. For benchmark purposes, we will evaluate the performance of using the global (weighted) clustering coefficient as a dissimilarity metric. In the same spirit we will also compare our proposal to spectral metrics, which are based on the comparison of the set of eigenvalues of the (weighted) adjacency or Laplacian matrices[33]. In the set of benchmark methods we will also evaluate WD-metric and Portrait Divergence. The former is defined as a linear combination of three distance terms[34, 35]: two of them compare, with appropriate metrics, the distributions of the (weighted) shortest-path lengths in the two graphs, while the third compares the distributions of the alpha-centralities in the two graphs and their complements. Portrait Divergence, on the other hand, defines a distance by comparing the ”portrait” of the two graphs, i.e., a matrix that encodes the distribution of the lengths of the shortest-path[36], and behaves very well in classifying binary networks[29]. It has been generalized to define a measure of dissimilarity between weighted networks (Ref.[36], Supplemental Material), although it has not been extensively tested in this regard. Finally, we note that the class of graphlet-based alignment-free methods [31, 32], which guarantees very high-level performance on binary networks [29, 30], does not yet have a generalization to weighted networks, as discussed in Ref.[37].
@@
The classification capabilities of the proposed ”weighted ego-distances” will be extensively evaluated on the testbed defined above. More precisely, the aim is to discriminate pairs of graphs originating from the same model from those originating from different models, with size and density (i.e., number of nodes and edges) acting as confounding factors. Note that the definition of ”model” includes the weighting scheme, so, for example, a dissimilarity measure should be able to recognize that two Barabási-Albert networks have different weighting schemes, even if they have exactly the same topology. The results of the experiments show that the proposed weighted ego-distances are able to outperform all distances used for comparison, and therefore should be considered, to date, the most advanced method for comparing undirected weighted networks. This statement is, of course, based on the set of benchmark measures and the composition of the network model pool used in the experimental setup: both, however, have been defined to the best of our knowledge.
The paper concludes with two application examples. In the first, a weighted ego-distance is used to compare the effects of different algorithms to filter weighted networks, i.e., to derive a ”backbone” by eliminating a large number of less significant edges while maintaining most of the total network weight[40]. Three filtering schemes are evaluated, namely the trivial hard threshold, the classical disparity filter[41], and the more recent Pólya filter[42]. Their application to two different datasets (the world trade network and the US airport network) reveals the superiority of the Pólya filter, in the sense that, at a given level of pruning, it is able to preserve the egonet (i.e., local) features of the original graph to a greater extent. The second application example concerns a temporal sequence of correlation networks, each obtained by analyzing the relationships between 3-month long time series referring to the stocks that define the S&P100 index. The pairwise comparison between the time-indexed correlation networks confirms a well-known result, namely that, given the greater coherence between the time series, the correlation graphs associated with financial crises have significantly different characteristics from those characterizing ordinary periods[43, 17, 44].
@@ Results
We evaluate the effectiveness of the ego-distances proposed above by generalizing the experimental setup defined in Refs. [31, 32, 30] with reference to unweighted networks. It is necessary to extend the above framework to the case of weighted networks for which, to the best of our knowledge, a standard benchmark pool of network models to be used for classification has never been defined to date.
@@
  • ER-x: Erdős-Rényi model with Uniform (ER-U), Random (ER-R), or Degree-dependent (ER-D) weighting.
  • BA-x: Barabási-Albert model with Uniform (BA-U), Random (BA-R), or Degree-dependent (BA-D) weighting.
  • GEO-x: geometric random graph model with Uniform (GEO-U), Random (GEO-R), or Degree-dependent (GEO-D) weighting.
  • YJBT: Yook-Jeong-Barabási-Tu model[38].
  • AK-x: Antal-Krapisvky model with Random (AK-R) or Exponential (AK-E) weighting[39].
@@
Figure 3: Example distributions of (normalized) weighted degree (left column), clustering coefficient (middle column), and egonet persistence (right column) for the $12$ synthetic network models ($N=1000$, $\rho=0.01$). Degree-homogeneous networks are in the top row, degree-heterogeneous networks in the bottom row (where a logarithmic x-axis is used to improve visibility).
@@
For each pair of networks, we calculate the eight ego-distances defined above (see Ego-distances). To have a comparison in terms of classification performance, we also calculate the following benchmark distances[29] (see Methods for details):
@@
  • $D_{WD}$ (WD-metric): It is based on a combined indicator that compares the distributions of the (weighted) shortest-path lengths of the two graphs, as well as the distributions of the alpha-centralities in the two graphs and their complements[34, 35].
  • $D_{PDiv}$ (Portrait Divergence): It is based on the comparison of the portrait matrices of the two networks, which encode the distribution of the (weighted) shortest-path lengths of the graphs[36].
@@
Table 1 summarizes, in terms of AUPR value, the results of the classification experiments described above. To combine the performance obtained by comparing graphs with different and with the same size/density, we also included the average of the two AUPR values (last column) as an overall indicator. In general, all ego-distances that use a combination of egonet features (i.e., $D_{SUM}$, $D_{cp}$, $D_{dc}$, $D_{dp}$, and $D_{dcp}$) perform better than the benchmark distances ($D_{C_{global}},D_{SP-W/L},{D_{WD}},D_{PDiv}$) in most cases. Partial exceptions are $D_{WD}$ (WD-metric) and $D_{PDiv}$ (Portrait Divergence), which perform similarly to ego-distances when comparing graphs of the same size/density, although their performance drops significantly more than ego-distances when considering mixed size/density. Overall, $D_{dcp}$, which exploits all the information available in the most extensive form, obtains the best performance on average. A viable and computationally simpler alternative is $D_{SUM}$, which achieves only slightly worse performance but is computationally lighter. Incidentally, we note that there is no clear evidence that one of the egonet features is less important than the other, to the point that it could be dropped: for example, $D_{c}$ performs the worst among the three 1-feature distances (see the last column avg), suggesting that the clustering coefficient might be less relevant, yet $D_{dc}$ – which does exploit the clustering coefficient – is the best among the 2-feature distances.
The Precision-Recall curves of Fig. 4 (left panel) show that the ego-distances which exploit all the available information ($D_{dcp}$ and $D_{SUM}$) achieve similar behavior to each other, and outperform the benchmark distances ($D_{PDiv}$ and to a larger extent $D_{C_{global}}$ and $D_{WD}$). $D_{dcp}$ provides the best trade-off at high Recall values, as evidenced by the largest F1 value slightly above $0.5$ (we recall that F1 is the harmonic mean of Precision and Recall). On the other hand, when the comparison is limited to networks with the same size/density, Fig. 4 (right panel) confirms that, as underlined in Table 1, $D_{dcp}$, $D_{SUM}$, and $D_{PDiv}$ achieve similar performance, slightly better than $D_{WD}$ at high Recall values, while $D_{C_{global}}$ follows with a large gap (further experiments on the dependence of the results on size and density are in the Supplementary Information file).
@@
distance all sizes/densities same size/density avg
@@
$D_{d},D_{c},D_{p}$: 1-feature ego-distances; $D_{cp},D_{dc},D_{dp}$: 2-feature ego-distances;
$D_{SUM},D_{dcp}$: 3-feature ego-distances; $D_{\text{C}_{global}}$: global clustering coefficient;
$D_{SP-W}$: spectral, weight matrix; $D_{SP-L}$: spectral, Laplacian matrix;
$D_{WD}$: WD-metric; $D_{PDiv}$: Portrait Divergence
Figure 4: Precision-Recall curves obtained by classifying the complete set of networks (12 models $\times$ 3 sizes $\times$ 3 densities $\times$ 10 replications = 1080 networks) with two of the ego-distances defined in the section Methods ($D_{SUM}$, $D_{dcp}$) and three of the benchmark distances ($D_{C_{global}}$, $D_{WD}$, $D_{PDiv}$) (the curves related to all other distances are omitted for readability). For each Recall-Precision point, the F1 value (i.e., the harmonic mean of Precision and Recall) is specified by the dotted contour line. Left: Performance is assessed by mixing sizes/densities; Right: Only networks with the same size/density are compared.
@@
Table 2: AUPR (Area Under the Precision/Recall curve) value for the classification of network models, for the ego-distances (top part of the table) and for the benchmark distances (bottom part) (we omit the distances $D_{d}$, $D_{c}$, $D_{p}$, and $D_{SP-W/L}$, whose performance is lower in most cases). The classification is performed separately for the network models with homogeneous or heterogeneous degree distribution. The best value in each column is in bold.
@@
A complementary analysis to the above is to group the network models based on their weighting scheme, rather than their degree distribution properties. This defines three groups, respectively with uniform (ER-/BA-/GEO-U), random (ER-/BA-/GEO-R, AK-R), or degree-dependent weighting (ER-/BA-/GEO-D, YJBT) – note that AK-E (with exponential weighting) is excluded because it has no other models to compare to. The results are in Table 3. Models with uniform weights (namely binary networks) are classified almost perfectly by ego-distances, despite mixed size/density acting as a confounder. Performance deteriorates slightly when considering random and degree-dependent weightings, although ego-distances remain significantly better. It is worth mentioning that much of the performance degradation reported in the second column (random weighting) is due to the inclusion of the AK-R model, which is only slightly different from BA-R (see Methods for details on the definition of the two models). If we remove AK-R from the pool and leave only ER-/BA-/GEO-R, the AUPR increases sharply for all measures (with the benchmark measures still dominated by ego-distances) reaching, e.g., $D_{dcp}=0.999$ or $D_{SUM}=0.993$. The same result occurs when removing YJBT from the third pool, given its similarity to BA-D (the two models have exactly the same topology and a slightly different weighting rule). Overall, this confirms that when considering a pool of networks with similar weighting scheme, the topological structure (not surprisingly) plays a crucial role for effective classification.
Table 3: AUPR (Area Under the Precision/Recall curve) value for the classification of network models, for the ego-distances (top part of the table) and for the benchmark distances (bottom part) (we omit the distances $D_{d}$, $D_{c}$, $D_{p}$, and $D_{SP-W/L}$, whose performance is lower in most cases). The classification is performed separately for the network models with uniform, random, or degree-dependent weighting. All sizes/densities are included in the experiment. The best value in each column is in bold.
uniform weighting random weighting degree-dependent weighting
distance ER-/BA-/GEO-U ER-/BA-/GEO-R, AK-R ER-/BA-/GEO-D, YJBT
$D_{SUM}$ 0.993 0.639 0.686
$D_{cp}$ 0.827 0.559 0.643
$D_{dc}$ 0.999 0.631 0.669
$D_{dp}$ 0.993 0.664 0.704
$D_{dcp}$ 0.999 0.683 0.730
$D_{\text{C}_{global}}$ 0.774 0.550 0.463
$D_{WD}$ 0.581 0.408 0.567
$D_{PDiv}$ 0.515 0.348 0.544
To summarize the results reported above, and to provide guidance on which of the proposed ego-distances to adopt, we firstly note that AUPR should be the primary selection criterion, as it represents the average precision over all possible recall values. When all network models are included in the classification problem (Table 1), it is not surprising that AUPR highlights the superiority of measures that exploit all available information: $D_{dcp}$ and $D_{SUM}$ have the best average performances, and the latter is a valid alternative to the former having a lower computational load (see Methods). Note that, in the ”all sizes/density” case, $D_{dcp}$ remains the best measure, with $D_{SUM}$ also achieving a very good performance, while in the ”same size/density” case the two measures, although not at the top, follow with a very small gap. The same excellent behavior of these two measures consistently emerges from the experiments in Tables 2 and 3, where the network models are split based on their degree distribution or weighting characteristics, respectively. In all cases, $D_{dcp}$ and $D_{SUM}$ provide the best performance or, in the worst case, a value close to the optimal one.
@@
The results are summarized in Fig. 5. As expected, for both networks and all filters used, the distance increases monotonically as more and more weight is removed. However, for any given value of the removed weight, the distance from the Hard threshold graph is greater than the distance from the graphs obtained by using the nontrivial filters (Disparity and Pólya), denoting that the former tends to perturb the original features to a larger extent. Furthermore, the Pólya filter produces backbone graphs much more similar to the original one than the Disparity filter, when the weight reduction is small. This is consistent with the properties of the Pólya filter, which has been shown to generate more heterogeneous backbones than those produced by other methods, i.e., more diverse than those obtained with ”naive” thresholding[42]. We note that, at the weight reduction level $R^{f}$ where the two curves become almost coincident (the two methods approximately coincide when the Pólya filter parameter reaches $a=1$, as theoretically discussed[42]), more than 80% of links are removed in both study cases, while maintaining at least 90% of the total weight. Therefore, the greater similarity with the original graph obtained from the Pólya filter is particularly appreciable, since it denotes that, for the same pruning effort, this scheme is able to preserve the egonet features of the original graph to a greater extent.
@@
Let us now use ego-distances to analyze data describing the stocks defining the S&P100 index, which includes most of the largest US companies. We start from the daily closing values $y_{i}^{\tau}$ of all stocks $i$ for the 10-year period January 2007 – December 2016 (Fig. 6(a)) and use them to calculate the daily returns $y_{i}^{\tau+1}/y_{i}^{\tau}$, which are then standardized to obtain time series $u_{i}^{\tau}$ with zero mean and unit variance. The 10-year time span is then divided into quarters $t=1,2,\ldots,40$ (that is, January 2006-March 2006, April 2006-June 2006, etc.). For each quarter, a correlation network is defined whose $100$ nodes correspond to the stocks. The nodes $(i,j)$ are connected with a weight given by $w_{ij}^{t}=(\rho_{ij}^{t}+1)/2$, where $\rho_{ij}^{t}$ is the Pearson correlation between the daily time series $(u_{i}^{\tau},u_{j}^{\tau})$ for the set of days $\tau$ of the quarter $t$. Thus $w_{ij}^{t}$ ranges from 00 (perfect anti-correlation) to $1$ (perfect correlation), with larger values denoting stronger agreement.
@@ Discussion
It is worth noting that the feature set could in principle be extended in many ways: weighted degree (strength), clustering coefficient, and persistence were chosen because they are egonet features, that is, for a node $i$ they can be computed only with knowledge of the properties of $i$ and its nearest neighbors. This simplifies the definition and implementation of the method and keeps the computational load at an acceptable level. It obviously involves a loss of information about the global graph structure, as for any embedding procedure — but designing a network comparison method boils down to finding precisely a good trade-off between simplicity (which implies loss of information) and effectiveness. We believe that our proposal produces a satisfactory balance, as witnessed by its good overall performance. Nonetheless, the proposed scheme could naturally be extended in many ways: while keeping the node-centric flavor, one could add one or more centrality indices (e.g., betweenness) to the three egonet features. Or one could go a step further and consider two-node feature distributions, as (weighted) pairwise similarity indices. Both extensions, however, would come at the cost of much higher computational requirements.
As a matter of fact, a significant limitation to the widespread use of the ego-distances could be their computational cost, which scales unfavorably with network size (see Methods) and therefore limits their applicability to small/medium sized graphs. It should be noted, however, that in principle the computation of egonet features can be fully parallelized, potentially allowing for a considerable reduction in computational requirements. Alternatively (or in combination), one could infer egonet distributions from only a sample of nodes, rather than from the entire network. This last approach, potentially promising for the analysis of large-scale networks, requires the use of appropriate graph sampling techniques[49]. We leave the in-depth evaluation of the effect of sampling on the performance of ego-distances to future research.